# Retro: Reusing teacher projection head for efficient embedding distillation on Lightweight Models via Self-supervised Learning

### v1

@@ Abstract
Self-supervised learning (SSL) is gaining attention for its ability to learn effective representations with large amounts of unlabeled data. Lightweight models can be distilled from larger self-supervised pre-trained models using contrastive and consistency constraints. Still, the different sizes of the projection heads make it challenging for students to mimic the teacher’s embedding accurately. We propose Retro, which reuses the teacher’s projection head for students, and our experimental results demonstrate significant improvements over the state-of-the-art on all lightweight models. For instance, when training EfficientNet-B0 using ResNet-50/101/152 as teachers, our approach improves the linear result on ImageNet to $66.9\%$, $69.3\%$, and $69.8\%$, respectively, with significantly fewer parameters.
keywords:
Self-supervised learning, light-weight model, contrastive learning.
@@ 1 Introduction
Deep learning has achieved remarkable success in various visual tasks, such as image classification, object detection, and semantic segmentation, thanks to the availability of large-scale annotated datasets. However, acquiring labeled data is time-consuming and expensive, making it crucial to explore better ways to utilize unlabeled data. Self-supervised learning (SSL) has emerged as an effective method to learn useful representations on unlabeled data, resulting in an outstanding performance on downstream tasks [11, 18, 7, 21, 3, 4, 12, 14].
Despite its effectiveness, most SSL methods require large networks, and the performance deteriorates when the model size is reduced. To address this issue, [9] proposed SEED, a self-supervised representation distillation method that distills the knowledge of larger pre-trained models into lightweight models via self-supervised learning. Similarly, CompRess [16] mimics the similarity score distribution between a teacher and a student over a dynamically maintained queue. [10] suggests incorporating consistency constraints between teacher and student embeddings to alleviate the Distilling Bottleneck problem via DisCo. BINGO [27] aims to transfer the relationship learned by the teacher to the student by leveraging a set of similar samples constructed by the teacher and grouped within a bag.
@@
Despite achieving state-of-the-art results across multiple tasks with high performance, some concerns remain. First, [10] discovered that expanding the dimension of the hidden layer in the MLP (projection head) could alleviate the Distilling Bottleneck problem. However, this approach is trivial since determining the size of the dimension and how large it should be remains unanswered. Second, because the student is lightweight with limited capability, it is challenging to accurately mimic the teacher from the encoder to the projection head. For example, in the DisCo study [10], they expanded the dimension to $2048$, which is the projection head dimension of ResNet-50/101/152. Consequently, this approach is equivalent to increasing the capability of mimicking the teacher, resulting in improved performance. However, when using ResNet-50$\times$2 with a dimension of $8192$ as a teacher, the performance on MobileNet-v3-Large and EfficientNet-B1 drops significantly and is inferior to the previous method [9]. Moreover, the feature distributions of the teacher and student models are statistically different and cannot be directly compared in practice, even if their dimensions are the same. Therefore, the optimal dimension for the projection head and how to efficiently distill the teacher embedding remain unanswered questions.
@@ 2.1 Self-supervised Learning
Recent studies have demonstrated the efficacy of contrastive-based techniques in self-supervised representation learning, where different perspectives of the same input are encouraged to be closer in feature space [3, 4, 5, 6, 12, 14]. For instance, SimCLR [3, 4] has proved that using strong data augmentation, larger batch sizes of negative samples, and including a projection head (MLP) after global average pooling can boost self-supervised learning. However, the performance of SimCLR is dependent on very large batch sizes and may not be feasible in real-world scenarios.
MoCo [6, 14], on the other hand, uses a memory bank to maintain consistent representations of negative samples. It considers contrastive learning as a look-up dictionary, enabling it to achieve superior performance without large batch sizes, making it more practical. BYOL [12] introduces a predictor to one branch of the network to prevent trivial solutions and break the symmetry. DINO [2] applies contrastive learning to vision transformers with self-distillation intuition.
@@
Figure 2: Comparison with existing self-supervised distillers. $x$ is the input image. The orange arrow indicates the knowledge transfer direction. Both 2a CompRess [16] and 2b SEED [9] transfer the knowledge of the similarity between a sample and a negative memory bank. 2c DisCo [10] constrains the last embedding of the student to be consistent with that of the teacher. 2d Our Retro improves DisCo by reusing the teacher projection head for the student, which has a higher capability to generate generalized embedding. ’Adt.’ indicates the adapter layer.
@@ 2.2 Knowledge Distillation
The idea of knowledge distillation was first proposed by [15], which transfers knowledge from a large teacher to a smaller student by minimizing the Kulback-Leibler (KL) divergence between the outputs of the two models. Attention Transfer (AT) is introduced to transfer the spatial attention of the teacher to the student by minimizing the mean squared error (MSE) between the feature maps of the two models. This method guides the student to focus on relevant regions of the input image, improving its performance on small datasets.
FitNets [23] is another method of knowledge distillation that transfers knowledge from the intermediate layers of a deep and thin teacher to a deeper but thinner student. The intermediate layers learned by the teacher are treated as hints, and the student is trained to mimic them using mean squared error loss. Relation Knowledge Distillation (RKD) [20] is a method that transfers the mutual relationship between the samples in a batch from the teacher to the student. RKD uses distance-wise and angle-wise distillation loss to transfer the relationship between the samples to the student.
@@ 3.1 Self-supervised Learning and Knowledge Distillation
In recent years, there has been a growing interest in combining knowledge distillation and self-supervised learning to improve the learning process. Some recent works, such as CRD [25] and SSKD [26], have used self-supervision as an auxiliary task to enhance knowledge distillation in fully supervised settings by transferring relationships between different modalities or mimicking transformed data and self-supervision tasks.
On the other hand, CompRess [16] and SEED [9] have focused on improving self-supervised visual representation learning on small models by incorporating knowledge distillation. They leverage the memory bank of MoCo [14] to maintain the consistency of the student’s distribution with that of the teacher. Meanwhile, DisCo [10] proposes to align the final embedding of the lightweight student with that of the teacher, exploiting the student’s learning ability to maximize knowledge. They also increase the dimension of the student’s projection head to better mimic the teacher’s embedding.
Figure 3: The pipeline of the proposed Retro technique. Two different data augmentation techniques first transform a single image into two views. A self-supervised pre-trained teacher is added in addition to the original contrastive SSL component, and the final embeddings generated by the learnable student and the frozen teacher must be consistent for each view. The contrastive is conducted with both views, and the projection head is frozen and reused in the student model.
@@ 3.2.1 Contrastive Learning Based SSL
To map the high-dimensional representations to a lower-dimensional embedding, a projection head $g\left(\cdot\right)$, which is a non-linear MLP, is used. Specifically, $g$ takes the representation $z$ as input and maps it to an embedding $E$ as $E=g\left(z\right)=g\circ f\left(x\right)$. These embeddings are then used to estimate the similarity in contrastive learning. The projection head is crucial to the success of self-supervised learning, as demonstrated in prior works such as MoCo [14] and SimCLR [3]. The encoder can be any network architecture, such as ResNet or EfficientNet.
@@ 3.2.2 DisCo
In DisCo, the input $x$ is transformed into two views $v$ and $v^{\prime}$ using two different augmentation strategies in each iteration. The views are fed into both the student encoder $f_{s}$ and the teacher encoder $f_{t}$ to create four representations $z_{s}$, $z^{\prime}_{s}$, $z_{t}$, and $z^{\prime}_{t}$. These representations are then projected using two different projection heads $g_{s}$ and $g_{t}$ to produce low-dimensional embeddings $E_{s}$, $E^{\prime}_{s}$, $E_{t}$, and $E^{\prime}_{t}$, respectively. The same process is also applied with the mean student, resulting in representations $z_{m}$, $z^{\prime}_{m}$ and embeddings $E_{m}$, $E^{\prime}_{m}$. The embeddings are then used to compute the contrastive learning loss using InfoNCE loss [19], similar to MoCo [14], as follows:
$\mathcal{L}_{\text{con}}=-\log\frac{\exp\left(\mathbf{q}\cdot\mathbf{k}^{+}/\tau\right)}{\sum_{i=0}^{K}\exp\left(\mathbf{q}\cdot\mathbf{k}_{i}/\tau\right)},$ (1)
@@
$\begin{gathered}\mathcal{L}_{\text{dis}}=\|E_{s}-E_{t}\|^{2}_{2}+\|E^{\prime}_{s}-E^{\prime}_{t}\|^{2}_{2}\end{gathered}$ (2)
@@ 3.3 Retro
The overall framework of Retro is illustrated in Figure 3. Retro comprises a lightweight student $s\left(\cdot\right)$, a mean student $m\left(\cdot\right)$, and a pre-trained frozen teacher $t\left(\cdot\right)$, which is similar to the DisCo framework [10]. However, unlike DisCo, we propose that the pre-trained teacher projection head can be used directly for the students since it contains the most valuable knowledge. Therefore, the objective is to train the student encoder to align the representation with the teacher projection head, instead of learning to mimic the teacher’s behavior. In addition, we leverage the power of the multi-view strategy by inputting both views into the mean student and enforcing the similarity constraint on pairs of embeddings between the student and the mean student.
@@ 3.4 Loss function and parameter update process
We follow BYOL [12] to symmetrize the contrastive loss in Retro as follows:
@@
Here, $q$ and $q^{\prime}$ are the embeddings from the student, while $k$ and $k^{\prime}$ are the embeddings from the mean student. We use two different memory banks for the two different views $v$ and $v^{\prime}$, respectively.
The overall loss function of Retro is formulated as follows:
$\mathcal{L}=\mathcal{L}_{\text{dis}}+\gamma\mathcal{L}_{\text{con}}$ (4)
@@ 4.1 Implementation Details
We first pre-train the self-supervised teacher models on the ImageNet dataset [24], which contains $1.3$ million training images and $50,000$ validation images with $1,000$ categories, using the MoCo-V2 [14] framework. We use ResNet as the backbone for the teacher models with different widths/depths, such as ResNet-50 (22.4M), ResNet-101 (40.5M), ResNet-152 (55.4M), and ResNet-50$\times$2 (94M). We pre-train ResNet-50/101 using the MoCo-V2 framework for $200$ epochs, ResNet-152 for $800$ epochs, while ResNet-50$\times$2 is pre-trained using the SwAV [1] method for $400$ epochs.
We evaluate our Retro method on five lightweight networks as students, including EfficientNet-B0 (4.0M), EfficientNet-B1 (6.4M), MobileNet-v3-Large (4.2M), ResNet-18 (10.7M), and ResNet-34 (20.4M). We use the same learning rate and optimizer as MoCo-V2 and train all student models for $200$ epochs. During distillation, the teacher’s and student’s projection heads are frozen for Retro, while SEED, DisCo, and BINGO only freeze the teacher. As a result, Retro has fewer trainable parameters and simpler training objectives. Note that SEED trains the models using a SwAV pre-trained teacher for $400$ epochs, and BINGO uses CutMix regularization [28] and more positive samples (5×5$\times$5 ×) during training, resulting in a higher benchmark score. Additionally, BINGO is not an end-to-end framework.
We later fine-tune the self-supervised distillation models for linear evaluation on ImageNet for $100$ epochs. We set the initial learning rate to $3$ for EfficientNet-B0/B1 and MobileNet-v3-Large, and $30$ for ResNet-18/34. The learning rate is scheduled to decrease by a factor of $10$ at $60$ and $80$ epochs, and we use SGD as the optimizer. We follow the other hyperparameters strictly as in MoCo-V2 [14].
@@ 4.2 Linear Evaluation
SEED [9] R-50 (67.4$\lozenge$ 61.3 82.7 61.4 83.1 55.2 80.3 57.6 81.8 58.5 82.6
DisCo [10] R-50 (67.4$\lozenge$ 66.5 87.6 66.6 87.5 64.4 86.2 60.6 83.7 62.5 85.4
BINGO [27] R-50 (67.4$\lozenge$ - - - - - - 61.4 84.3 63.5 85.7
@@
SEED [9] R-101 (70.3) 63.0 83.8 63.4 84.6 59.9 83.5 58.9 82.5 61.6 84.9
DisCo [10] R-101 (69.1$\lozenge$ 68.9 88.9 69.0 89.1 65.7 86.7 62.3 85.1 64.4 86.5
@@
SEED [9] R-152 (74.2) 65.3 86.0 67.3 86.9 61.4 84.6 59.5 83.3 62.7 85.8
DisCo [10] R-152 (74.1$\lozenge$ 67.8 87.0 73.1 91.2 63.7 84.9 65.5 86.7 68.1 88.6
BINGO [27] R-152 (74.1$\lozenge$ - - - - - - 65.9 87.1 69.1 88.9
@@
SEED [9] R-50$\times$2 (77.3$\dagger$ 67.6 87.4 68.0 87.6 68.2 88.2 63.0 84.9 65.7 86.8
DisCo [10] R-50$\times$2 (77.3$\dagger$ 69.1 88.9 64.0 84.6 58.9 81.4 65.2 86.8 67.6 88.6
BINGO [27] R-50$\times$2 (77.3$\dagger$ - - - - - - 65.5 87.0 68.9 89.0
@@
4.3 Semi-supervised Linear Evaluation
We also evaluate our method in semi-supervised scenarios, following previous methodologies. We use $1\%$ and $10\%$ sampled subsets of the ImageNet training data (i.e., $13$ and $128$ samples per class, respectively) to fine-tune the student models. As shown in Table 2, our Retro approach consistently outperforms the baseline under any quantity of labeled data. Notably, our method achieves these results while strictly following the settings from SEED [9] and DisCo [10], whereas BINGO uses a higher learning rate ($10$) for the classifier layer.
@@ 4.3 Semi-supervised Linear Evaluation
SEED[9] R-50 (67.4) 39.1 50.2
DisCo[10] R-50 (67.4) 39.2 50.1
BINGO[27] R-50 (67.4) 42.8 57.5
@@
SEED[9] R-152 (74.1) 44.3 54.8
DisCo[10] R-152 (74.1) 47.1 54.7
BINGO[27] R-152 (74.1) 50.3 61.2

### v3

# National Cancer Center

@@ Abstract
Self-supervised learning (SSL) is gaining attention for its ability to learn effective representations with large amounts of unlabeled data. Lightweight models can be distilled from larger self-supervised pre-trained models using contrastive and consistency constraints, but the different sizes of the projection heads make it challenging for students to accurately mimic the teacher’s embedding. We propose Retro, which reuses the teacher’s projection head for students, and our experimental results demonstrate significant improvements over the state-of-the-art on all lightweight models. For instance, when training EfficientNet-B0 using ResNet-50/101/152 as teachers, our approach improves the linear result on ImageNet to $66.9\%$, $69.3\%$, and $69.8\%$, respectively, with significantly fewer parameters.
@@ 1 Introduction
Deep learning has achieved remarkable success in various visual tasks, such as image classification, object detection, and semantic segmentation, thanks to the availability of large-scale annotated datasets. However, acquiring labeled data is time-consuming and expensive, making it crucial to explore better ways to utilize unlabeled data. Self-supervised learning (SSL) has emerged as an effective method to learn useful representations on unlabeled data, resulting in an outstanding performance on downstream tasks [Gidaris et al. (2018), Noroozi and Favaro (2016), Doersch et al. (2015), Pathak et al. (2016), Chen et al. (2020a), Chen et al. (2020b), Grill et al. (2020), He et al. (2020)].
Despite its effectiveness, most SSL methods require large networks, and the performance deteriorates when the model size is reduced. To address this issue, (Fang et al., 2021) proposed SEED, a self-supervised representation distillation method that distills the knowledge of larger pre-trained models into lightweight models via self-supervised learning. Similarly, CompRess Koohpayegani et al. (2020) mimics the similarity score distribution between a teacher and a student over a dynamically maintained queue. (Gao et al., 2022) suggests incorporating consistency constraints between teacher and student embeddings to alleviate the Distilling Bottleneck problem via DisCo. BINGO Xu et al. (2021) aims to transfer the relationship learned by the teacher to the student by leveraging a set of similar samples constructed by the teacher and grouped within a bag.
@@
Despite achieving state-of-the-art results across multiple tasks with high performance, some concerns remain. First, Gao et al. (2022) discovered that expanding the dimension of the hidden layer in the MLP (projection head) could alleviate the Distilling Bottleneck problem. However, this approach is trivial since determining the size of the dimension and how large it should be remains unanswered. Second, because the student is lightweight with limited capability, it is challenging to accurately mimic the teacher from the encoder to the projection head. For example, in the DisCo study Gao et al. (2022), they expanded the dimension to $2048$, which is the projection head dimension of ResNet-50/101/152. Consequently, this approach is equivalent to increasing the capability of mimicking the teacher, resulting in improved performance. However, when using ResNet-50$\times$2 with a dimension of $8192$ as a teacher, the performance on MobileNet-v3-Large and EfficientNet-B1 drops significantly and is inferior to the previous method Fang et al. (2021). Moreover, the feature distributions of the teacher and student models are statistically different and cannot be directly compared in practice, even if their dimensions are the same. Therefore, the optimal dimension for the projection head and how to efficiently distill the teacher embedding remain unanswered questions.
@@ 2.1 Self-supervised Learning
Recent studies have demonstrated the efficacy of contrastive-based techniques in self-supervised representation learning, where different perspectives of the same input are encouraged to be closer in feature space Chen et al. (2020a, b); Chen and He (2021); Chen et al. (2020c); Grill et al. (2020); He et al. (2020). For instance, SimCLR Chen et al. (2020a, b) has proved that using strong data augmentation, larger batch sizes of negative samples, and including a projection head (MLP) after global average pooling can boost self-supervised learning. However, the performance of SimCLR is dependent on very large batch sizes and may not be feasible in real-world scenarios.
MoCo Chen et al. (2020c); He et al. (2020), on the other hand, uses a memory bank to maintain consistent representations of negative samples. It considers contrastive learning as a look-up dictionary, enabling it to achieve superior performance without large batch sizes, making it more practical. BYOL Grill et al. (2020) introduces a predictor to one branch of the network to prevent trivial solutions and break the symmetry. DINO Caron et al. (2021) applies contrastive learning to vision transformers with self-distillation intuition.
@@
Figure 2: Comparison with existing self-supervised distillers. $x$ is the input image. The orange arrow indicates the knowledge transfer direction. Both 2(a) CompRess Koohpayegani et al. (2020) and 2(b) SEED Fang et al. (2021) transfer the knowledge of the similarity between a sample and a negative memory bank. 2(c) DisCo Gao et al. (2022) constrains the last embedding of the student to be consistent with that of the teacher. 2(d) Our Retro improves DisCo by reusing the teacher projection head for the student, which has a higher capability to generate generalized embedding. ’Adt.’ indicates the adapter layer.
@@ 2.2 Knowledge Distillation
The idea of knowledge distillation was first proposed by (Hinton et al., 2015), which transfers knowledge from a large teacher to a smaller student by minimizing the Kulback-Leibler (KL) divergence between the outputs of the two models. Attention Transfer (AT) is introduced to transfer the spatial attention of the teacher to the student by minimizing the mean squared error (MSE) between the feature maps of the two models. This method guides the student to focus on relevant regions of the input image, improving its performance on small datasets.=
FitNets Romero et al. (2014) is another method of knowledge distillation that transfers knowledge from the intermediate layers of a deep and thin teacher to a deeper but thinner student. The intermediate layers learned by the teacher are treated as hints, and the student is trained to mimic them using mean squared error loss. Relation Knowledge Distillation (RKD) Park et al. (2019) is a method that transfers the mutual relationship between the samples in a batch from the teacher to the student. RKD uses distance-wise and angle-wise distillation loss to transfer the relationship between the samples to the student.
@@ 3.1 Self-supervised Learning and Knowledge Distillation
In recent years, there has been a growing interest in combining knowledge distillation and self-supervised learning to improve the learning process. Some recent works, such as CRD Tian et al. (2019) and SSKD Xu et al. (2020), have used self-supervision as an auxiliary task to enhance knowledge distillation in fully supervised settings by transferring relationships between different modalities or mimicking transformed data and self-supervision tasks.
On the other hand, CompRess Koohpayegani et al. (2020) and SEED Fang et al. (2021) have focused on improving self-supervised visual representation learning on small models by incorporating knowledge distillation. They leverage the memory bank of MoCo He et al. (2020) to maintain the consistency of the student’s distribution with that of the teacher. Meanwhile, DisCo Gao et al. (2022) proposes to align the final embedding of the lightweight student with that of the teacher, exploiting the student’s learning ability to maximize knowledge. They also increase the dimension of the student’s projection head to better mimic the teacher’s embedding.
@@ 3.2.1 Contrastive Learning Based SSL
To map the high-dimensional representations to a lower-dimensional embedding, a projection head $g\left(\cdot\right)$, which is a non-linear MLP, is used. Specifically, $g$ takes the representation $z$ as input and maps it to an embedding $E$ as $E=g\left(z\right)=g\circ f\left(x\right)$. These embeddings are then used to estimate the similarity in contrastive learning. The projection head is crucial to the success of self-supervised learning, as demonstrated in prior works such as MoCo He et al. (2020) and SimCLR Chen et al. (2020a). The encoder can be any network architecture, such as ResNet or EfficientNet.
@@
Figure 3: The pipeline of the proposed Retro technique. Two different data augmentation techniques first transform a single image into two views. A self-supervised pre-trained teacher is added in addition to the original contrastive SSL component, and the final embeddings generated by the learnable student and the frozen teacher must be consistent for each view.
@@ 3.2.2 DisCo
In DisCo, the input $x$ is transformed into two views $v$ and $v^{\prime}$ using two different augmentation strategies in each iteration. The views are fed into both the student encoder $f_{s}$ and the teacher encoder $f_{t}$ to create four representations $z_{s}$, $z^{\prime}_{s}$, $z_{t}$, and $z^{\prime}_{t}$. These representations are then projected using two different projection heads $g_{s}$ and $g_{t}$ to produce low-dimensional embeddings $E_{s}$, $E^{\prime}_{s}$, $E_{t}$, and $E^{\prime}_{t}$, respectively. The same process is also applied with the mean student, resulting in representations $z_{m}$, $z^{\prime}_{m}$ and embeddings $E_{m}$, $E^{\prime}_{m}$. The embeddings are then used to compute the contrastive learning loss using InfoNCE loss Oord et al. (2018), similar to MoCo He et al. (2020), as follows:
$\small\mathcal{L}_{\text{con}}=-\log\frac{\exp\left(\mathbf{q}\cdot\mathbf{k}^{+}/\tau\right)}{\sum_{i=0}^{K}\exp\left(\mathbf{q}\cdot\mathbf{k}_{i}/\tau\right)},$ (1)
@@
$\small\begin{gathered}\mathcal{L}_{\text{dis}}=\|E_{s}-E_{t}\|^{2}_{2}+\|E^{\prime}_{s}-E^{\prime}_{t}\|^{2}_{2}\end{gathered}$ (2)
@@ 3.3 Retro
The overall framework of Retro is illustrated in Figure 3. Retro comprises a lightweight student $s\left(\cdot\right)$, a mean student $m\left(\cdot\right)$, and a pre-trained frozen teacher $t\left(\cdot\right)$, which is similar to the DisCo framework Gao et al. (2022). However, unlike DisCo, we propose that the pre-trained teacher projection head can be used directly for the students since it contains the most valuable knowledge. Therefore, the objective is to train the student encoder to align the representation with the teacher projection head, instead of learning to mimic the teacher’s behavior. In addition, we leverage the power of the multi-view strategy by inputting both views into the mean student and enforcing the similarity constraint on pairs of embeddings between the student and the mean student.
@@ 3.4 Loss function and parameter update process
We follow BYOL Grill et al. (2020) to symmetrize the contrastive loss in Retro as follows:
@@
Here, $q$ and $q^{\prime}$ are the embeddings from the student, while $k$ and $k^{\prime}$ are the embeddings from the mean student. We use two different memory banks for the two different views $v$ and $v^{\prime}$, respectively. The overall loss function of Retro is formulated as follows:
$\small\mathcal{L}=\mathcal{L}_{\text{dis}}+\gamma\mathcal{L}_{\text{con}}$ (4)
@@ 4.1 Implementation Details
We first pre-train the self-supervised teacher models on the ImageNet dataset Russakovsky et al. (2015), which contains $1.3$ million training images and $50,000$ validation images with $1,000$ categories, using the MoCo-V2 He et al. (2020) framework. We use ResNet as the backbone for the teacher models with different widths/depths, such as ResNet-50 (22.4M), ResNet-101 (40.5M), ResNet-152 (55.4M), and ResNet-50$\times$2 (94M). We pre-train ResNet-50/101 using the MoCo-V2 framework for $200$ epochs, ResNet-152 for $800$ epochs, while ResNet-50$\times$2 is pre-trained using the SwAV Caron et al. (2020) method for $400$ epochs.
We evaluate our Retro method on five lightweight networks as students, including EfficientNet-B0 (4.0M), EfficientNet-B1 (6.4M), MobileNet-v3-Large (4.2M), ResNet-18 (10.7M), and ResNet-34 (20.4M). We use the same learning rate and optimizer as MoCo-V2 and train all student models for $200$ epochs. During distillation, the teacher’s and student’s projection heads are frozen for Retro, while SEED, DisCo, and BINGO only freeze the teacher. As a result, Retro has fewer trainable parameters and simpler training objectives. Note that SEED trains the models using a SwAV pre-trained teacher for $400$ epochs, and BINGO uses CutMix regularization Yun et al. (2019) and more positive samples (5×5$\times$5 ×) during training, resulting in a higher benchmark score. Additionally, BINGO is not an end-to-end framework.
We later fine-tune the self-supervised distillation models for linear evaluation on ImageNet for $100$ epochs. We set the initial learning rate to $3$ for EfficientNet-B0/B1 and MobileNet-v3-Large, and $30$ for ResNet-18/34. The learning rate is scheduled to decrease by a factor of $10$ at $60$ and $80$ epochs, and we use SGD as the optimizer. We follow the other hyperparameters strictly as in MoCo-V2 He et al. (2020).
@@ 4.2 Linear Evaluation
SEED Fang et al. (2021) R-50 (67.4$\lozenge$ 61.3 82.7 61.4 83.1 55.2 80.3 57.6 81.8 58.5 82.6
DisCo Gao et al. (2022) R-50 (67.4$\lozenge$ 66.5 87.6 66.6 87.5 64.4 86.2 60.6 83.7 62.5 85.4
BINGO Xu et al. (2021) R-50 (67.4$\lozenge$ - - - - - - 61.4 84.3 63.5 85.7
@@
SEED Fang et al. (2021) R-101 (70.3) 63.0 83.8 63.4 84.6 59.9 83.5 58.9 82.5 61.6 84.9
DisCo Gao et al. (2022) R-101 (69.1$\lozenge$ 68.9 88.9 69.0 89.1 65.7 86.7 62.3 85.1 64.4 86.5
@@
SEED Fang et al. (2021) R-152 (74.2) 65.3 86.0 67.3 86.9 61.4 84.6 59.5 83.3 62.7 85.8
DisCo Gao et al. (2022) R-152 (74.1$\lozenge$ 67.8 87.0 73.1 91.2 63.7 84.9 65.5 86.7 68.1 88.6
BINGO Xu et al. (2021) R-152 (74.1$\lozenge$ - - - - - - 65.9 87.1 69.1 88.9
@@
SEED Fang et al. (2021) R-50$\times$2 (77.3$\dagger$ 67.6 87.4 68.0 87.6 68.2 88.2 63.0 84.9 65.7 86.8
DisCo Gao et al. (2022) R-50$\times$2 (77.3$\dagger$ 69.1 88.9 64.0 84.6 58.9 81.4 65.2 86.8 67.6 88.6
BINGO Xu et al. (2021) R-50$\times$2 (77.3$\dagger$ - - - - - - 65.5 87.0 68.9 89.0
@@
SEEDFang et al. (2021) R-50 (67.4) 39.1 50.2
DisCoGao et al. (2022) R-50 (67.4) 39.2 50.1
BINGOXu et al. (2021) R-50 (67.4) 42.8 57.5
@@
SEEDFang et al. (2021) R-152 (74.1) 44.3 54.8
DisCoGao et al. (2022) R-152 (74.1) 47.1 54.7
BINGOXu et al. (2021) R-152 (74.1) 50.3 61.2
@@
4.3 Semi-supervised Linear Evaluation
We also evaluate our method in semi-supervised scenarios, following previous methodologies. We use $1\%$ and $10\%$ sampled subsets of the ImageNet training data (i.e., $13$ and $128$ samples per class, respectively) to fine-tune the student models. As shown in Table 2, our Retro approach consistently outperforms the baseline under any quantity of labeled data. Notably, our method achieves these results while strictly following the settings from SEED Fang et al. (2021) and DisCo Gao et al. (2022), whereas BINGO uses a higher learning rate ($10$) for the classifier layer.